Two hours after Harlan had left his house for his out-of-state business trip, his wife Maggie noticed that he had forgotten his briefcase. Knowing that the briefcase contained all of his important papers, she immediately hopped into her car and began to chase after him at a speed of 65 mph. If Harlan drove 55 mph the entire time, how many hours did it take Maggie to catch him?
Found 2 solutions by solver91311, ankor@dixie-net.com:
Answer by solver91311(24713)



How far does Harlan travel in the first two hours at 55 mph?

How long does it take to travel that distance at 10 miles per hour (the difference in their speeds)?


John

My calculator said it, I believe it, that settles it


Answer by ankor@dixie-net.com(22740)

Two hours after Harlan had left his house for his out-of-state business trip, his wife Maggie noticed that he had forgotten his briefcase.
Knowing that the briefcase contained all of his important papers, she immediately hopped into her car and began to chase after him at a speed of 65 mph.
If Harlan drove 55 mph the entire time, how many hours did it take Maggie to catch him?
:
let t = travel time of M
then
(t+2) = travel time of H, who started 2 hrs earlier
:
When M catches H, they will have traveled the same distance
Write a distance equation, dist = speed * time
65t = 55(t+2)
65t = 55t + 110
65t - 55t = 110
10t = 110
t = 110/10
t = 11 hrs !
:
Check this, find the actual distance each traveled
M: 65(11) = 715 mi
H: 55(13) = 715 mi